What are the concentrations of all the solute species in 1.05 M lactic acid, HC3H5O2? What is the pH of the solution? This acid has a Ka = 1.4 × 10-4.
Answer #1

Lactic acid is a weak acid.

According to Ostwald's theory of weak electroytes,

[H3O+] = [C3H5O2-] = sqrt.(Ka*C) = sqrt(1.4*10-4*1.05) = 0.0121 M

[HC3H5O2] = 1.05 - 0.0121 = 1.0379 M

pH = - Log[H3O+] = - Log(0.0121) = 1.92